8. You buy an eight-year bond that has a 6% current yield and a 6% coupon rate (coupons will be paid annually). The face value is $1000. In one year, the yield-to-maturity of this bond has dropped to 5%. What is the bond’s holding-period return? ____%
Answer #1

Current yield = Coupon / Price

Coupon = 0.06 * 1000 = 60

Price = Coupon /Current yield

Price = 60 / 0.06

Price = $1,000

When price is equal to face value, yield to maturity equal to will be equal to coupon rate. Therefore, Yield to maturity will be 6%.

Price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Price = 60 * [1 - 1 / (1 + 0.05)7] / 0.05 + 1000 / (1 + 0.05)7

Price = 60 * [1 - 0.71068] / 0.05 + 710.68133

Price = 60 * 5.78637 + 710.68133

Price = $1,057.86373

holding-period return = [(1,057.86373 + 60 - 1,000) / 1,000] * 100

holding-period return = 11.79%